### v2

# Next-to-leading-order QCD corrections to nucleon Dirac form factors

Abstract
The leading-order perturbative QCD (pQCD) predictions to nucleon electromagnetic form factors were first made in late 70s. Nearly half a century later, we accomplish the calculation of the next-to-leading-order (NLO) QCD corrections to proton and neutron’s Dirac form factors at large momentum transfer, to the leading-twist accuracy in collinear factorization. The NLO hard-scattering kernel is computed in the Krankl and Manashov renormalization scheme. The effect of NLO perturbative corrections turns out to be positive and substantial. Confronting our state-of-the-art pQCD predictions with the available data for nucleon Dirac form factors, in both space-like and time-like regions, imposes a stringent test on the validity of the parameterized forms of nucleon leading-twist light-cone-distribution amplitudes (LCDAs), which have been predicted by a class of QCD sum rules-based models and lattice QCD simulation.
Introduction. As the composite particles made of strongly-coupled, relativistic quarks and gluons, proton and neutron are the building block of the atomic nuclei, consequently, of our planet and all its inhabitants, and even more, of our visible universe. Understanding the internal structure of nucleons from quantum chromodynamics (QCD), the underlying theory of strong interaction, is the fundamental challenge faced by the contemporary hadron and nuclear physics Gross:2022hyw .
Nucleon electromagnetic form factors (EMFFs), which gauge the distributions of the electric charge and magnetization inside the nucleon, are the fundamental probes to the internal structure of nucleon Gross:2022hyw . The nucleon EMFFs can be directly accessed from the $ep$ elastic scattering experiment, which are encoded in the following nucleon transition matrix element:
$\langle P^{\prime}|j^{\mu}_{\rm em}|P\rangle=\bar{N}\left[F_{1}(Q^{2})\gamma^{\mu}-F_{2}(Q^{2})\frac{\sigma^{\mu\nu}q_{\nu}}{2m_{N}}\right]N,$ (1)
with the electromagnetic current $j^{\mu}_{\rm em}=\sum_{f}e_{f}\bar{\psi}_{f}\gamma^{\mu}\psi_{f}$ ($e_{u}={2\over 3}$, $e_{d}=e_{s}=-{1\over 3})$. $N$, $m_{N}$ denote the Dirac spinor and mass of the nucleon, respectively. $q^{\mu}\equiv P^{\mu}-P^{\prime\mu}$, and $Q^{2}\equiv-q^{2}\geq 0$ characterizes the squared momentum transfer. The scalar functions $F_{1}$, $F_{2}$ are referred to as the (helicity-conserving) Dirac form factor, and the (helicity-flipped) Pauli form factor, respectively.
After Hofstader’s pioneering work in mid-50s Hofstadter:1955ae ; Mcallister:1956ng , intensive experimental efforts have been continuously devoted to measuring the nucleon EMFFs $F^{p,n}_{1,2}$ over the past seven decades (for comprehensive experimental review, see Gao:2003ag ; Hyde:2004gef ; Perdrisat:2006hj ; Punjabi:2015bba ). In early 90s SLAC has already measured the space-like proton EMFFs at $Q^{2}=31.3\;{\rm GeV}^{2}$ Sill:1992qw . In the prospective electron-ion colliders EIC and EicC, the momentum transfer for nucleon EMFFs may be further extended to $Q^{2}\sim 40-50\;{\rm GeV}^{2}$ Schmookler:2022gxw . Over the past two decades, the high luminosity $e^{+}e^{-}$ colliders has also started to explore the nucleon EMFFs in the time-like region Antonelli:1998fv ; BES:2005lpy ; CLEO:2005tiu ; Seth:2012nn ; BaBar:2013ves ; BaBar:2013ukx ; BESIII:2021tbq .
On the theoretical side, enormous intellectual endeavors have also been invested to unveiling the QCD dynamics beneath the nucleon EMFFs. At small and moderate $Q^{2}$, the measured nucleon EMFFs have been confronted with the theoretical predictions made from various nonperturbative approaches, e.g., vector meson dominance model Iachello:1972nu ; Gari:1984ia ; Lomon:2006xb ; Brodsky:2003gs ; Dubnickova:1992ii ; Chen:2023oqs , constitute quark model Chung:1991st ; Faessler:2005gd ; deMelo:2008rj , cloudy bag model Thomas:1981vc ; Betz:1983dy ; Song:1992ei , Dyson-Schwinger equation Bashir:2012fs ; Cloet:2013gva ; Segovia:2014aza , light-front Hamiltonian Mondal:2019jdg , chiral perturbation theory Schindler:2005ke ; Bauer:2012pv , light-cone sum rule Braun:2001tj , lattice QCD Collins:2011mk ; Alexandrou:2011db ; QCDSF:2017ssq ; Capitani:2015sba ; Alexandrou:2018sjm ; Jang:2019jkn , as well as dispersive analysis Mergell:1995bf ; Belushkin:2006qa .
Thanks to asymptotic freedom, the nucleon EMFFs with large momentum transfer are generally believed to be adequately accounted by perturbative QCD (pQCD). Within the collinear factorization formalism Lepage:1979zb ; Lepage:1979za ; Lepage:1980fj ; Efremov:1978rn ; Efremov:1979qk ; Duncan:1979ny ; Duncan:1979hi ; Chernyak:1977fk , they can be expressed as the perturbatively-calculable hard-scattering kernel convoluted with nonperturbative light-cone distribution amplitudes (LCDAs) of nucleon. Helicity selection rule implies that, at asymptotically large $Q^{2}$, $F_{1}$ scales with $1/Q^{4}$, and $F_{2}$ scales with $1/Q^{6}$ Brodsky:1974vy , modulo logarithms of $Q^{2}$ Belitsky:2002kj . Hence nucleon Dirac form factor dominates Pauli form factor in such a limit.
The lowest-order (LO) hard-scattering kernel for $F_{1}$ was known in the late 70s, shortly after the advent of QCD Lepage:1979za ; Chernyak:1980dj ; Lepage:1980fj ; Chernyak:1977fk . Curiously, notwithstanding nearly half a centaury has elapsed, the next-to-leading order (NLO) perturbative correction to this fundamental observable of nucleon remains unknown until today, obviously due to severe technical obstacles  [^6] Note the status of pQCD predictions for nucleons EMFFs lags far behind that for pion EMFFs, where the one-loop QCD corrections were known in early 80s Field:1981wx ; Dittes:1981aw ; Sarmadi:1982yg ; Braaten:1987yy , and even the two-loop corrections have also become available very recently Chen:2023byr .. The theoretical stagnation for such a basic observable over a very long time is rather unusual, which should be contrasted with the $ep$ deep inelastic scattering, where the three-loop QCD corrections became available about two decades ago Vermaseren:2005qc .
It is the very goal of this work to fill this overdue gap, by accomplishing the calculation of the NLO perturbative correction to the nucleon Dirac form factors, and conducting a comprehensive phenomenological study.
Collinear factorization for nucleon Dirac form factor. At large momentum transfer, the $F_{1}$ can be factorized as the following convolution integrals:
$Q^{4}F_{1}(Q^{2})$ $=$ $\int\!\!\!\!\int[dx][dy]\,\Phi_{3}^{*}(y,\mu_{F})\,\Phi_{3}(x,\mu_{F})$ (2)
×$\times\;T_{H}\left(x,y,Q^{2},\mu_{R}^{2},\mu_{F}^{2}\right)+{\cal O}(1/Q^{2}),$
with the integration measure $[dx]\equiv dx_{1}dx_{2}dx_{3}\delta(1-\sum_{i}x_{i})$. $T_{H}(x,y)$ signifies the hard-scattering kernel, with $\mu_{R}$, $\mu_{F}$ being the renormalization and factorization scales. $\Phi_{3}(x)$ represents the twist-3 nucleon LCDA, $i.e.$, the probability amplitude of finding the three valence quarks inside nucleon carrying the momentum fractions $x_{i}$ ($i=1,2,3$). This factorized structure has been picturised in a cartoon in Fig. 1.
Figure 1: Pictorial illustration of the nucleon EMFF in collinear factorization approach. For concreteness, we show one typical LO and NLO Feynman diagrams for partonic reactions.
For definiteness let us specialize to the proton. The leading-twist-3 LCDA of the proton with helicity $h={1\over 2}$ assumes the following operator definition Chernyak:1983ej ; Braun:2000kw :
[…]
with $P^{2}=m_{N}^{2}$. $i,j,k=1,\cdots,3$ denote the color indices, and $q_{\pm}\equiv{1\over 2}(1\pm\gamma_{5})q$ represents the right and left-handed components of the massless quark field, and $u^{\uparrow}_{p}(P)$ signifies the Dirac spinor of $h={1\over 2}$ proton in relativistic normalization. ${\cal C}\equiv i\gamma^{2}\gamma^{0}$ represents the charge conjugation matrix. $n^{\mu}$ is a light-like vector. For simplicity we have suppressed those light-like gauge links whose role is to ensure gauge invariance. Renormalizing the nonlocal baryonic operator in (3), one recovers the celebrated Efremov-Radyushkin-Brodsky-Lepage (ERBL) evolution equation Lepage:1980fj ; Efremov:1979qk :
$\frac{d\Phi_{3}(x,\mu_{F})}{d\ln\mu^{2}_{F}}=V(x,y)\mathop{\otimes}\limits_{y}\Phi_{3}(y,\mu_{F}),$ (4)
where $\mathop{\otimes}\limits_{y}$ is a shorthand for the convolution integration over $[dy]$.
Perturbative matching in KM scheme. We first sketch how to achieve the desired factorization formula (2). With $Q^{2}\gg\Lambda^{2}_{\rm QCD}$, the transition matrix element in (1) in $d$-dimensional spacetime can be factorized as
$\langle\Pi^{\mu}(Q)\rangle$$\equiv$ $\langle P^{\prime}|j_{\rm em}^{\mu}|P\rangle=\langle P^{\prime}|{\bm{[}}\bar{\cal Q}_{\alpha\beta\gamma}(y){\bm{]}}^{\overline{\rm MS}}|0\rangle$
$\;\mathop{\otimes}_{y}\;{\cal T}_{\alpha\beta\gamma;\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x,y)\;\mathop{\otimes}_{x}\;\langle 0|{\bm{[}}{\cal Q}_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x){\bm{]}}^{\overline{\rm MS}}|P\rangle,$
where $\alpha,\beta,\gamma,\cdots$ denote the spinor indices affiliated with each quark line, ${\cal T}$ signifies the perturbative coefficient function. The three-quark operator ${\cal Q}_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x)$ is defined via the following Fourier transform Anikin:2013aka :
$\widetilde{\cal Q}_{\alpha\beta\gamma}(\{a_{i}\})=\int[dx]e^{-iP_{n}\sum a_{i}x_{i}}{\cal Q}_{\alpha\beta\gamma}(x),$ (6)
where $\widetilde{\cal Q}_{\alpha\beta\gamma}(\{a_{i}\})$ represents the following non-local three-quark operators with light-like separations:
$\widetilde{\cal Q}_{\alpha\beta\gamma}(\{a_{i}\})=\epsilon^{ijk}u^{i}_{\alpha}(a_{1}n)u^{j}_{\beta}(a_{2}n)d^{k}_{\gamma}(a_{3}n),$ (7)
with the light-like gauge links tacitly inserted.
The ${\overline{\rm MS}}$-renormalized three-quark operators in (Next-to-leading-order QCD corrections to nucleon Dirac form factors) is connected with the bare one through
${\bm{[}}{\cal Q}_{\alpha\beta\gamma}(x){\bm{]}}^{\overline{\rm MS}}={\cal Z}_{\alpha\beta\gamma}^{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x,x^{\prime})\mathop{\otimes}_{x^{\prime}}{\cal Q}_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x^{\prime}).$ (8)
It is advantageous to adopt the renormalization scheme originally proposed by Krankl and Manashov Krankl:2011gch (KM scheme henceforth). The KM scheme has several remarkable virtues. It guarantees the vanishing of evanescent operators Dugan:1990df in $d=4$ dimensions so that one only needs to work on the four-dimensional operator basis. The renormalization procedure in KM scheme preserves Fierz identities between the renormalized operators Anikin:2013aka . The self-consistency of the KM scheme has been checked through the three-loop accuracy Gracey:2012gx . Moreover, this scheme has also been utilized by the lattice QCD practitioners to decipher the baryon LCDA Bali:2015ykx ; RQCD:2019hps ; Bali:2024oxg .
The ${\cal Z}$ factor in (8) can be parameterized in KM scheme as
${\cal Z}_{\alpha\beta\gamma}^{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x,x^{\prime})$ $=$ $\delta_{\alpha\beta\gamma}^{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}\delta(x-x^{\prime})$ (9)
$+\sum_{s=1}{1\over\epsilon^{s}}\,a^{s}_{lmn}\,\gamma^{(l)}_{\alpha\alpha^{\prime}}\otimes\gamma^{(m)}_{\beta\beta^{\prime}}\otimes\gamma^{(n)}_{\gamma\gamma^{\prime}}$
where $\gamma^{(n)}$ is the shorthand for the totally anti-symmetrized product of $n$ $\gamma$-matrices, $\gamma^{(n)}_{\mu_{1}\cdots\mu_{n}}\equiv\gamma^{[\mu_{1}}\gamma^{\mu_{2}}\cdots\gamma^{\mu_{n}]}$. Note that all the Lorentz indices of $\gamma$-matrices are contracted between different quark lines, where there is only one independent way to contract all indices in ${\cal Z}$. For reader’s convenience, the one-loop expression of ${\cal Z}$ is explicitly given in Appendix A.
The perturbative coefficient function $\cal T$ in (Next-to-leading-order QCD corrections to nucleon Dirac form factors) can be decomposed into
${\cal T}_{\alpha\beta\gamma;\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x,y)=T_{lmn}(x,y)\gamma^{(l)}_{\alpha\alpha^{\prime}}\otimes\gamma^{(m)}_{\beta\beta^{\prime}}\otimes\gamma^{(n)}_{\gamma\gamma^{\prime}}.$ (10)
Note that the hard kernel $T$ is a scalar function, and Lorentz index $\mu$ is affiliated with one of the $\gamma^{(l)}$, $\gamma^{(m)}$ and $\gamma^{(n)}$ matrices.
The central task now is to deduce the coefficient function $T_{lmn}(x,y)$. Since it is insensitive to the long-distance physics, one can replace the incoming and outgoing nucleons by a fictitious nucleon, viz., the free, collinear three-quark states $|q_{1}(u_{1}P)q_{2}(u_{2}P)q_{3}(u_{3}P)\rangle$ and $|q_{1}(v_{1}P)q_{2}(v_{2}P)q_{3}(v_{3}P)\rangle$. As a great simplification of KM scheme, the quark flavor becomes irrelevant in determining ${\cal Z}$ and $T$ Krankl:2011gch .
For a fictitious nucleon, the left-side in (Next-to-leading-order QCD corrections to nucleon Dirac form factors) can be computed in perturbation theory, which can also be expanded in the basis spanned by the totally anti-symmetrized product of $\gamma$-matrices:
$\langle\Pi^{\mu}(Q)\rangle={\cal A}_{lmn}(u,v)\,\bar{u}_{\alpha\beta\gamma}\left[\gamma^{(l)}_{\alpha\alpha^{\prime}}\otimes\gamma^{(m)}_{\beta\beta^{\prime}}\otimes\gamma^{(n)}_{\gamma\gamma^{\prime}}\right]u_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}},$
where $u_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}\equiv u_{\alpha^{\prime}}(u_{1}P)u_{\beta^{\prime}}(u_{2}P)u_{\gamma^{\prime}}(u_{3}P)$.
Analogously, we also evaluate the nonlocal matrix elements in the right-side of (Next-to-leading-order QCD corrections to nucleon Dirac form factors) perturbatively. After dropping all the scaleless integrals, we obtain
$\langle 0|{\bm{[}}{\cal Q}_{\alpha\beta\gamma}(x){\bm{]}}^{\overline{\rm MS}}|P\rangle={\cal Z}_{\alpha\beta\gamma}^{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x,u)\,u_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}.$ (12)
Finally we arrive at the following matching equation to deduce $T$:
${\cal A}_{lmn}(u,v)\gamma^{(l)}_{\alpha\alpha^{\prime}}\otimes\gamma^{(m)}_{\beta\beta^{\prime}}\otimes\gamma^{(n)}_{\gamma\gamma^{\prime}}=\bar{\cal Z}_{\alpha\beta\gamma}^{\alpha_{2}\beta_{2}\gamma_{2}}(y,v)$ (13)
$\mathop{\otimes}_{y}\left[T_{l^{\prime}m^{\prime}n^{\prime}}(x,y)\gamma^{(l^{\prime})}_{\alpha_{2}\alpha_{1}}\otimes\gamma^{(m^{\prime})}_{\beta_{2}\beta_{1}}\otimes\gamma^{(n^{\prime})}_{\gamma_{2}\gamma_{1}}\right]\mathop{\otimes}_{x}{\cal Z}_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}^{\alpha_{1}\beta_{1}\gamma_{1}}(x,u).$
Note that one has to reexpand the products of $\gamma$-marices in the right-hand side, and finally cast them in the basis of the totally anti-symmetrized products of $\gamma$-matrices.
Sketch of the calculation. We start with the partonic reaction $\gamma^{*}+q_{1}(u_{1}P)q_{2}(u_{2}P)q_{3}(u_{3}P)\to q_{1}(v_{1}P^{\prime})q_{2}(v_{2}P^{\prime})q_{3}(v_{3}P^{\prime})$, We use HepLib Feng:2021kha and FeynArts Hahn:2000kx to generate the Feynman diagrams and the corresponding amplitudes. There are about 48 tree-level diagrams, and 2040 one-loop diagrams.
As is evident in Fig. 1, the amplitude for each Feynman diagram entails the following form:
$\bar{u}_{\alpha\beta\gamma}\left[\Gamma^{(1)}_{\alpha\alpha^{\prime}}\otimes\Gamma^{(2)}_{\beta\beta^{\prime}}\otimes\Gamma^{(3)}_{\gamma\gamma^{\prime}}\right]u_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}},$ (14)
After moving $P\!\!\!\!/$ to the rightmost and $P\!\!\!\!/\,^{\prime}$ to the leftmost repeatedly using the anti-communication relation, and with the aid of Dirac equation, each $\Gamma^{(i)}$ can be expressed in terms of the products of a string of Dirac $\gamma$ matrices, which can be further cast into the linear combination of the totally anti-symmetrized products of $\gamma$ matrices.
The aforementioned operations can be automated by the package HepLib Feng:2021kha . We then use Apart Feng:2012iq and FIRE Smirnov:2019qkx for partial fraction and IBP reduction. We end up with about 100 one-loop two-point and three-point master integrals, which are computed analytically by the differential equation method Kotikov:1990kg ; Remiddi:1997ny . Upon renormalizing the QCD coupling in $\overline{\rm MS}$ scheme, we obtain a UV finite expression of the NLO QCD amplitude ${\cal A}_{lmn}$ in (11), which however still possess uncancelled single infrared pole. From the matching equation (13), we find that, hearteningly, the IR poles in the NLO amplitude are exactly cancelled by the corresponding IR poles in the ${\cal Z}$ factor, so that the short-distance coefficient function $T_{lmn}(x,y)$ is indeed IR finite in a pointwise manner. Therefore it is verified that, the collinear factorization does hold at one-loop order in the baryonic sector.
To recover the form of the standard collinear factorization formula in (2), we substitute the following expression Braun:1999te ; Anikin:2013aka
$\langle 0|{\bm{[}}{\cal Q}_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x){\bm{]}}^{\overline{\rm MS}}|P\rangle=\frac{1}{4}\Big{\{}{\cal V}(P\!\!\!\!/{\cal C})_{\alpha\beta}(\gamma_{5}N)_{\gamma}$ (15)
$\hskip 28.45274pt+{\cal A}(P\!\!\!\!/\gamma_{5}{\cal C})_{\alpha\beta}N_{\gamma}+{\cal T}(i\sigma_{\mu P}{\cal C})_{\alpha\beta}(\gamma_{\mu}\gamma_{5}N)_{\gamma}\Big{\}}$
into (Next-to-leading-order QCD corrections to nucleon Dirac form factors), and use the following relation between ${\cal V}(x),\,{\cal A}(x),\,{\cal T}(x)$ and $\Phi_{3}(x)$:
$\Phi_{3}(x_{1},x_{2},x_{3})={\cal V}(x_{1},x_{2},x_{3})-{\cal A}(x_{1},x_{2},x_{3}),$ (16a)
${\cal T}(x_{1},x_{2},x_{3})=\frac{1}{2}\left[\Phi_{3}(x_{1},x_{3},x_{2})+\Phi_{3}(x_{2},x_{3},x_{1})\right],$
and obtain
$Q^{4}\langle\Pi^{\mu}(Q)\rangle=\Phi_{3}(x)\mathop{\otimes}_{x}T_{H}(x,y)\mathop{\otimes}_{y}\Phi_{3}^{*}(y)\;\bar{N}\gamma^{\mu}N.$ (17)
One then extracts the intended Dirac form factor through NLO accuracy as indicated in (2).
Master formula for nucleon Dirac form factor at NLO. The leading-twist-3 nucleon LCDA is conveniently expanded in the orthogonal polynomials basis Braun:2008ia :
$\Phi_{3}(x,\mu_{F})$ $=$ $f_{\rm N}(\mu_{F})\!\!\sum_{0\leq i\leq j}\!\!a_{ij}(\mu_{F})\phi_{ij}(x)$ (18)
$\phi_{ij}(x)$ $=$ $120x_{1}x_{2}x_{3}\mathcal{P}_{ij}(x),$ (19)
with $f_{\rm N}$ denotes the nucleon decay constant. The first few orthogonal polynomials ${\cal P}_{ij}(x)$ have been tabulated in Braun:2008ia ; Braun:2014wpa (e.g, (12) in Braun:2014wpa ). All the nonperturbative dynamics is encoded in the expansion coefficients $a_{ij}(\mu_{F})$.
For phenomenological usage, it is advantageous to recast (2) from the convolution form into an algebraic one:
$Q^{4}F_{1}(Q^{2})=f_{\rm N}^{2}(\mu_{F}){\sum_{ijkl}}a_{ij}(\mu_{F}){\cal H}^{ijkl}a_{kl}(\mu_{F}),$ (20)
with ${\cal H}^{ijkl}$ defined by
${\cal H}^{ijkl}=\phi_{ij}(x)\mathop{\otimes}\limits_{x}T_{H}(x,y,\frac{\mu_{R}^{2}}{Q^{2}},\frac{\mu_{F}^{2}}{Q^{2}})\mathop{\otimes}\limits_{y}\phi_{kl}(y).$ (21)
The short-distance coefficients ${\cal H}^{ijkl}$ can be calculated in perturbation theory. Through NLO accuracy, they can be parameterized as
${\cal H}^{ijkl}$ $=$ $\frac{1600\pi^{2}\alpha_{s}^{2}}{3}\Big{\{}e_{u}\,c^{ijkl}_{0}\left[1+{\alpha_{s}\over\pi}\left(c^{ijkl}_{1}L_{\mu}+c^{ijkl}_{2}\right)\right]$ (22)
$+e_{d}\,d^{ijkl}_{0}\left[1+{\alpha_{s}\over\pi}(d^{ijkl}_{1}L_{\mu}+d^{ijkl}_{2})\right]+{\cal O}(\alpha_{s}^{2})\Big{\}},$
with $L_{\mu}\equiv\ln\mu^{2}/Q^{2}$. For simplicity we have set $\mu_{R}=\mu_{F}=\mu$. We find $c^{ijkl}_{1}=d^{ijkl}_{1}$, and ${\cal H}^{ijkl}={\cal H}^{klij}$ originating from $T_{H}(x,y)=T_{H}(y,x)$. For a given $ijkl$, the coefficients $c^{ijkl}_{012}$ and $d^{ijkl}_{012}$ can be analytically inferred. For concreteness, we just list their values for the case of $0000$:
${\cal H}^{0000}$ $=$ $\frac{1600\alpha_{s}^{2}\pi^{2}}{3}\Big{\{}e_{u}\left[1+\frac{\alpha_{s}}{\pi}\left(\frac{29}{6}L_{\mu}-92\zeta_{5}\right.\right.$ (23)
$\left.\left.+\frac{3719}{30}\zeta_{3}-\frac{1897}{60}\right)\right]+2e_{d}\left[1+\frac{\alpha_{s}}{\pi}\left(\frac{29}{6}L_{\mu}+\frac{3}{2}\zeta_{5}\right.\right.$
$\left.\left.+\frac{209}{120}\zeta_{3}+\frac{1741}{120}\right)\right]+{\cal O}(\alpha_{s}^{2})\Big{\}},$
${ijkl}$ $c^{ijkl}_{0}$ $c^{ijkl}_{2}$ $d^{ijkl}_{0}$ $d^{ijkl}_{2}$ $c^{ijkl}_{1}=d^{ijkl}_{1}$
$0000$ 1.0000 22.001 2.0000 18.157 4.8333
$1000$ 8.1667 25.030 -8.1667 25.030 5.3889
$1010$ 168.78 29.453 174.22 28.937 5.9444
$1100$ 3.5000 29.699 -3.5000 18.566 5.5000
$1110$ 38.111 30.441 -38.111 30.441 6.0556
$1111$ 59.889 27.617 21.778 26.486 6.1667
$2000$ 21.350 26.952 8.0500 25.082 5.7222
$2010$ 89.833 32.057 -89.833 32.057 6.2778
$2011$ 93.100 32.942 9.8000 38.805 6.3889
$2020$ 234.71 35.077 54.880 34.146 6.6111
$2100$ 8.7500 28.194 -8.7500 28.194 5.9444
$2110$ 159.25 33.344 159.25 33.569 6.5000
$2111$ 44.917 33.321 -44.917 33.321 6.6111
$2120$ 84.525 36.543 -84.525 36.543 6.8333
$2121$ 159.25 38.254 159.25 38.232 7.0556
$2200$ 3.2000 27.963 0.10000 47.505 6.0000
$2210$ 12.250 32.193 -12.250 32.193 6.5556
$2211$ 7.1167 40.486 2.6833 35.971 6.6667
$2220$ 30.345 36.568 4.9350 37.529 6.8889
$2221$ 7.1750 39.572 -7.1750 39.572 7.1111
$2222$ 7.7900 35.348 1.7200 35.302 7.1667
Table 1: Numerical values of $c_{i}$ and $d_{i}$ affiliated with various ${\cal H}^{ijkl}$ with $0\leq\{ijkl\}\leq 2$.
Conducting the four-fold convolution integrations to accurately inferring ${\cal H}^{ijkl}$ turns out to be a great challenge, largely due to severe numerical instability caused by many fake singularities in $T_{H}^{(1)}(x,y)$, which are located either at the boundaries or in the interior of the integration domain. We conduct the four-fold integration from two different routes: one is to utilize the numerical integrator Cubature cubature implementing multiple-precision floating-point around 100 digits, and the other is to accomplish the integration analytically with the assistance of the package PolyLogTools Duhr:2019tlz . The perfect agreement is found between the analytic and numerical approaches. In Table 1 we enumerate the values for $c_{i}$ and $d_{i}$ associated with a few lowest-lying $ijkl$  [^7] We notice that, a PhD thesis by Knödlseder in 2015 was dedicated to calculating the NLO QCD correction to nucleon Dirac form factor at leading twist Knodlseder:2015vmu . Though the analytic result of the one-loop amplitude was obtained in Knodlseder:2015vmu , the IR-finite hard-scattering kernel $T^{(1)}(x,y)$ has not been given in a pointwise manner. More importantly, since the author failed to accomplish the four-fold convolution integrations in a numerically stable fashion, no attempt of confronting the NLO pQCD predictions with the measured nucleon form factor data was made in Knodlseder:2015vmu ..
Equipped with Table 1, Equation (20) constitutes the master formula for yielding phenomenological predictions for proton Dirac form factor through the one-loop accuracy. In the isospin symmetric limit, the neutron Dirac form factor $F_{1}^{n}$ can be obtained from (20) by making the replacement $e_{u}\leftrightarrow e_{d}$ in (22). It is also straightforward to adapt this master formula from the space-like region to the time-like one, provided that one makes the replacement $L_{\mu}\to L_{\mu}+i\pi$ in (22), with $Q^{2}$ now indicating the squared invariant mass of the nucleon-antinucleon pair.
Phenomenological exploration. It is now the time to confront our state-of-the-art pQCD predictions with the available data. We take the large-$Q^{2}$ data points for spacelike proton Dirac form factor from SLAC 1986 Coward:1967au and SLAC 1993 Sill:1992qw   [^8] Note it is the magnetic Sachs form factor of nucleon, $G_{\rm M}(Q^{2})$, that is directly accessible in $ep$ scattering experiments. At large $Q^{2}$, it nevertheless becomes legitimate to assume $F_{1}(Q^{2})\approx G_{\rm M}(Q^{2})$.. We take the data points for time-like proton Dirac form factor from BES BES:2005lpy , CLEO CLEO:2005tiu , NU Seth:2012nn , E835 E835:1999mlt , E760 E760:1992rvj , LA ISR BaBar:2013ves and SA ISR BaBar:2013ukx . We discard many irrelevant data at small momentum transfer. In contrast, the quality of data and the covered range of $Q^{2}$ for neutron EMFFs is much less competitive with respect to that for proton. The data points for spacelike neutron Dirac form factor are taken from ROCK Rock:1982gf , Lung Lung:1992bu , while those for timelike neutron Dirac form factor are taken from FENICE Antonelli:1998fv and BES3 BESIII:2021tbq .
Figure 2: pQCD predictions v.s. data for $Q^{4}F^{p}_{1}(Q^{2})$ in the space-like region. The values of the expansion coefficients of proton LCDA are taken with eight different ansatzs. The green and red bands correspond to the LO and NLO pQCD predictions, which are obtained by sliding $\mu_{R}=\mu_{F}=\mu$ from $Q/2$ to $Q$. Since the pQCD predictions in the lower row are much smaller than the data, we have multiplied them with a magnifying factor (as specified inside the parenthesis in each plot) to render the theoretical bands visible.
Figure 3: Same as Fig. 2, except switching from space-like to time-like region.
Figure 4: Similar to Fig. 2, but with proton replaced with the neutron. Here pQCD predictions are confronted with the data for $Q^{4}F^{n}_{1}(Q^{2})$ in the space-like region.
Figure 5: Same as Fig. 4, except switching from space-like to time-like region.
As the key nonperturative input, our present knowledge about the nucleon LCDAs is far from satisfactory and confirmative. During the past decades, The low-lying expansion coefficients of nucleon LCDAs have been extensively investigated Chernyak:1984bm ; Chernyak:1987nv ; King:1986wi ; Stefanis:1992nw ; Braun:2006hz ; Anikin:2013aka ; Braun:2014wpa ; RQCD:2019hps . In our numerical analysis, we have taken eight different ansatzs of nucleon LCDAs as input. Six of them are obtained from the QCD sum rules, referred to as CZ Chernyak:1984bm , COZ Chernyak:1987nu , KS King:1986wi , SB Stefanis:1992nw , BLW Braun:2006hz and ABO Anikin:2013aka , respectively. Two of them are provided by the recent lattice QCD simulations, abbreviated as LAT 14 Braun:2014wpa and RQCD RQCD:2019hps , respectively.
The predicted values of the expansion coefficients of nucleon are usually given at the initial scale $\mu_{0}=1$ or 2 GeV, which are scattered in a wide range in literature. The one-loop renormalization group equations are invoked to evolve $f_{\rm N}$ and $a_{nk}(\mu)$ from $\mu_{0}$ to another scale, viz. $f_{\rm N}(\mu)=f_{\rm N}(\mu_{0})\left[\alpha_{s}(\mu)/\alpha_{s}(\mu_{0})\right]^{2/(3\beta_{0})}$ and $a_{ij}(\mu)=a_{ij}(\mu_{0})\left[\alpha_{s}(\mu)/\alpha_{s}(\mu_{0})\right]^{\gamma_{ij}/\beta_{0}}$. $\beta_{0}=11-{2\over 3}n_{f}$ with $n_{f}=3$ signifies the one-loop coefficient of the QCD $\beta$ function, while $\gamma_{ij}$ is the anomalous dimension associated with each expansion coefficient, whose explicit form can be found in Braun:2014wpa . The package FAPT Bakulev:2012sm is utilized to evaluate the running QCD coupling constant to two-loop accuracy.
From Fig. 2 through Fig. 5, we confront our state-of-the-art pQCD predictions with the available large-$Q^{2}$ data for proton and neutron’s Dirac form factors, in both space-like and time-like cases. The theoretical errors are estimated by varying $\mu_{R}=\mu_{F}=\mu$ from $Q/2$ to $Q$.
It has been widely known that the LO predictions in pQCD for proton Dirac form factor are generally much smaller than the large-$Q^{2}$ data, for virtually all the popular models of nucleon LCDAs Stefanis:1997zyh . To resolve this discrepancy, it has been proposed Chernyak:1984bm ; Gari:1986ue ; Stefanis:1987vr ; Gari:1986dr ; Ji:1986uh ; Stefanis:1992nw that the strong coupling constants should be evaluated at the exact gluon virtualities in each diagram, so that the LO prediction can be significantly enhanced to account for the data. The $\alpha_{s}$ in this recipe might be associated with a scale of order $\Lambda_{\rm QCD}$, evaluation of which requires some ad hoc treatment. Moreover, it is difficult to adapt this recipe to the NLO case.
As is evident from Fig. 2 through Fig. 5, the magnitudes of the NLO QCD correction are positive and significant. After including the NLO QCD correction, the pQCD predictions, within errors, can decently account for the large-$Q^{2}$ data for both proton and neutron’s Dirac form factors, in both spacelike and timelike regions, once we adopt the CZ Chernyak:1984bm , COZ Chernyak:1987nu , KS King:1986wi , SB Stefanis:1992nw parametrizations of nucleon LCDAs. This phenomenological success of NLO prediction implies that the frozen $\alpha_{s}$ trick is unnecessary and should be abandoned.
In sharp contrast, due to very small expansion coefficients predicted by BLW Braun:2006hz , ABO Anikin:2013aka , LAT14 Braun:2014wpa and RQCD RQCD:2019hps , the resulting pQCD predictions are far below the experimental data. From Fig. 2 through Fig. 5, one clearly sees that, even after including NLO corrections, the pQCD predictions based on these models are still order-of-magnitude smaller than the measured Dirac form factor of proton and neutron.
Since lattice QCD is universally believed to be the only model-independent nonperturbative approach to unravel the hadronic structure, the sheer failure of the parameterized nucleon LCDAs provided by LAT14 and RQCD when confronting with the data is rather disquieting. This nuisance is apparently due to too small values of the Appel moments determined by the lattice method. To settle this alarming discrepancy, it seems imperative to call for another independent lattice study based on a diametrically different algorithm. The large momentum effective theory (LaMET) Ji:2013dva is a viable option, which allows one to ascertain hadron’s parton distributions directly in the $x$ space on Euclidean lattice (for a recent review on LaMET, see Ji:2020ect ). Some theoretical investigations have recently been launched to infer light octet baryons’ LCDAs within the LaMET framework Deng:2023csv ; Han:2024ucv .
Summary and Outlook. In this work we report the first successful calculation of the NLO QCD corrections to the nucleon Dirac form factors. The celebrated collinear factorization formalism has been explicitly verified to hold true at one-loop order for hard-exclusive process involving baryons. Somewhat surprisingly, it takes nearly half a century to advance the LO prediction to the NLO level for such a basic observable related to nucleon internal structure.
It has long been known that the LO prediction in pQCD severely underestimates the measured nucleon Dirac form factor data. Historically some ad hoc frozen strong coupling recipe were advocated to reconcile this discrepancy. Remarkably, the effect of NLO perturbative corrections turns to be positive and significant. After including the NLO QCD correction, the pQCD predictions can decently describe the large-$Q^{2}$ data for both proton and neutron’s Dirac form factors, in both spacelike and timelike regions, provided that the parameterized forms of the nucleon LCDAs are taken from a class of QCD sum rules-based models, which renders the aforementioned frozen $\alpha_{s}$ trick unnecessary and obsolete. On the other hand, some other models and lattice QCD studies predict too small expansion coefficients of nucleon LCDAs, so that the resulting pQCD predictions, even after including NLO corrections, are still orders-of-magnitude smaller than the measured nucleon form factors.
In a sense, this work may herald the coming of a new phase of perturbative QCD, where NLO QCD corrections start to be systematically explored for a class of important hard exclusive reactions in baryon sector, exemplified by the nucleon Pauli (axial, gravitational) form factors, nucleon Compton scattering, $\gamma\gamma\to N\overline{N}$, quarkonium decay into $N\overline{N}$, and so on. It is conceivable that the effect of NLO QCD corrections could be significant in many occasions, including which will thereby facilitate a critical comparison between pQCD prediction and data.
Acknowledgments. We are grateful to Deshan Yang for valuable discussion. The work of L. B. C. is supported in part by NNSFC under Grant No. 12175048. The work of W. C. is supported by Guangdong Major Project of Basic and Applied Basic Research (No. 2020B0301030008). The work of F. F. is supported in part by NNSFC under Grant No. 12275353. The work of S. H. and Y. J. is supported in part by NNSFC under Grant No. 11925506.
Note added. After the first version of this paper was submitted to arXiv, there has appeared a similar work by Huang, Shi, Wang and Zhao [arXiv:2407.18724[hep-ph]], who have also computed the NLO QCD correction to nucleon Dirac form factor in the spacelike domain, yet within a specific evanescent-to-physical operator mixing scheme.
Appendix A One-loop renormalization factor of non-local three-quark operators in KM scheme
To the best of our knowledge, the explicit one-loop expression of the renormalization factor ${\cal Z}$ in (8), which is associated with the nonlocal three-quark opeartors with light-like separations, has never been given in literature. For reader’s convenience, here we explicit present its closed form in KM scheme Krankl:2011gch , accurate through the one-loop accuracy:
${\cal Z}$ $=$ $\delta(x_{1}-x^{\prime}_{1})\delta(x_{2}-x^{\prime}_{2})\Gamma_{000}-\frac{\alpha_{s}}{12\pi}\frac{1}{\epsilon}\Bigg{\{}18\delta(x_{1}-x^{\prime}_{1})\delta(x_{2}-x^{\prime}_{2})\Gamma_{000}$ (24)
$\left.+\Gamma_{000}\left\{\delta(x^{\prime}_{3}-x_{3})\left[\frac{4x_{2}}{x^{\prime}_{2}}\left[\frac{\theta(x_{1}-x^{\prime}_{1})}{x_{1}-x^{\prime}_{1}}\right]_{+}+\frac{4x_{1}}{x^{\prime}_{1}}\left[\frac{\theta(x_{2}-x^{\prime}_{2})}{x_{2}-x^{\prime}_{2}}\right]_{+}-\frac{4(x^{\prime}_{1}-x_{1})\theta(x_{2}-x^{\prime}_{2})}{x^{\prime}_{1}x^{\prime}_{2}}+\frac{4x_{2}}{x^{\prime}_{2}(1-x^{\prime}_{3})}\right]\right.\right.$
$+\delta(x^{\prime}_{2}-x_{2})\left[\frac{4x_{1}}{x^{\prime}_{1}(1-x^{\prime}_{2})}+\frac{4(x^{\prime}_{1}-x_{1})\theta(x_{1}-x^{\prime}_{1})}{x^{\prime}_{1}x^{\prime}_{3}}+\frac{4x_{1}}{x^{\prime}_{1}}\left[\frac{\theta(x_{3}-x^{\prime}_{3})}{x_{3}-x^{\prime}_{3}}\right]_{+}+\frac{4x_{3}}{x^{\prime}_{3}}\left[\frac{\theta(x_{1}-x^{\prime}_{1})}{x_{1}-x^{\prime}_{1}}\right]_{+}\right]$
$\left.+\delta(x^{\prime}_{1}-x_{1})\left[\frac{4x_{2}}{(1-x^{\prime}_{1})x^{\prime}_{2}}+\frac{4(x^{\prime}_{2}-x_{2})\theta(x_{2}-x^{\prime}_{2})}{x^{\prime}_{2}x^{\prime}_{3}}+\frac{4x_{2}}{x^{\prime}_{2}}\left[\frac{\theta(x_{3}-x^{\prime}_{3})}{x_{3}-x^{\prime}_{3}}\right]_{+}+\frac{4x_{3}}{x^{\prime}_{3}}\left[\frac{\theta(x_{2}-x^{\prime}_{2})}{x_{2}-x^{\prime}_{2}}\right]_{+}\right]\right\}$
$+\delta(x^{\prime}_{2}-x_{2})\Gamma_{202}\left[\frac{(x^{\prime}_{1}-x_{1})\theta(x_{1}-x^{\prime}_{1})}{x^{\prime}_{1}x^{\prime}_{3}}+\frac{x_{1}}{x^{\prime}_{1}(1-x^{\prime}_{2})}\right]+\delta(x^{\prime}_{1}-x_{1})\Gamma_{022}\left[\frac{(x^{\prime}_{2}-x_{2})\theta(x_{2}-x^{\prime}_{2})}{x^{\prime}_{2}x^{\prime}_{3}}+\frac{x_{2}}{(1-x^{\prime}_{1})x^{\prime}_{2}}\right]$
$+\delta(x^{\prime}_{3}-x_{3})\Gamma_{220}\left[\frac{x_{2}}{x^{\prime}_{2}(1-x^{\prime}_{3})}-\frac{(x^{\prime}_{1}-x_{1})\theta(x_{2}-x^{\prime}_{2})}{x^{\prime}_{1}x^{\prime}_{2}}\right]\Bigg{\}}+{\cal O}(\alpha_{s}^{2}),$
where the basis matrices $\Gamma_{ijk}$ are defined by
$\Gamma_{000}=I\otimes I\otimes I,\quad\Gamma_{220}=\gamma^{(2)}_{\mu\nu}\otimes\gamma^{(2)}_{\mu\nu}\otimes I,\quad\Gamma_{202}=\gamma^{(2)}_{\mu\nu}\otimes I\otimes\gamma^{(2)}_{\mu\nu},\quad\Gamma_{022}=I\otimes\gamma^{(2)}_{\mu\nu}\otimes\gamma^{(2)}_{\mu\nu}.$ (25)
The $+$-distribution is defined by
[…]
where $f(x_{i})$ signifies a test function.
Moreover, with the aid of the following identities,
$\frac{x_{2}}{x^{\prime}_{2}}\left[\frac{\theta(x_{1}-x^{\prime}_{1})}{x_{1}-x^{\prime}_{1}}\right]_{+}=\left[\frac{x_{2}}{x^{\prime}_{2}}\frac{\theta(x^{\prime}_{2}-x_{2})}{x^{\prime}_{2}-x_{2}}\right]_{+}-\delta(x_{1}-x^{\prime}_{1}),$ (27)
$\frac{x_{2}}{x^{\prime}_{2}(1-x^{\prime}_{3})}-\frac{(x^{\prime}_{1}-x_{1})\theta(x_{2}-x^{\prime}_{2})}{x^{\prime}_{1}x^{\prime}_{2}}=\left[\frac{x_{2}}{x^{\prime}_{2}}\theta(x^{\prime}_{2}-x_{2})+\frac{x_{1}}{x^{\prime}_{1}}\theta(x^{\prime}_{1}-x_{1})\right]\frac{1}{(1-x^{\prime}_{3})},$ (28)
one can obtain an alternative expression for ${\cal Z}$:
${\cal Z}=\delta(x_{1}-x^{\prime}_{1})\delta(x_{2}-x^{\prime}_{2})\Gamma_{000}-\frac{\alpha_{s}}{12\pi}\frac{1}{\epsilon}\Bigg{\{}-6\delta(x_{1}-x^{\prime}_{1})\delta(x_{2}-x^{\prime}_{2})\Gamma_{000}$
$+\Gamma_{000}\left\{\delta(x^{\prime}_{3}-x_{3})\left\{\left[\frac{4x_{2}}{x^{\prime}_{2}}\frac{\theta(x^{\prime}_{2}-x_{2})}{x^{\prime}_{2}-x_{2}}+\frac{4x_{1}}{x^{\prime}_{1}}\frac{\theta(x^{\prime}_{1}-x_{1})}{x^{\prime}_{1}-x_{1}}\right]_{+}+\left[\frac{4x_{2}}{x^{\prime}_{2}}\theta(x^{\prime}_{2}-x_{2})+\frac{4x_{1}}{x^{\prime}_{1}}\theta(x^{\prime}_{1}-x_{1})\right]\frac{1}{1-x^{\prime}_{3}}\right\}\right.$
$+\delta(x^{\prime}_{2}-x_{2})\left\{\left[\frac{4x_{3}}{x^{\prime}_{3}}\frac{\theta(x^{\prime}_{3}-x_{3})}{x^{\prime}_{3}-x_{3}}+\frac{4x_{1}}{x^{\prime}_{1}}\frac{\theta(x^{\prime}_{1}-x_{1})}{x^{\prime}_{1}-x_{1}}\right]_{+}+\left[\frac{4x_{3}}{x^{\prime}_{3}}\theta(x^{\prime}_{3}-x_{3})+\frac{4x_{1}}{x^{\prime}_{1}}\theta(x^{\prime}_{1}-x_{1})\right]\frac{1}{1-x^{\prime}_{2}}\right\}$
$+\delta(x^{\prime}_{1}-x_{1})\left\{\left[\frac{4x_{2}}{x^{\prime}_{2}}\frac{\theta(x^{\prime}_{2}-x_{2})}{x^{\prime}_{2}-x_{2}}+\frac{4x_{3}}{x^{\prime}_{3}}\frac{\theta(x^{\prime}_{3}-x_{3})}{x^{\prime}_{3}-x_{3}}\right]_{+}+\left[\frac{4x_{2}}{x^{\prime}_{2}}\theta(x^{\prime}_{2}-x_{2})+\frac{4x_{3}}{x^{\prime}_{3}}\theta(x^{\prime}_{3}-x_{3})\right]\frac{1}{1-x^{\prime}_{1}}\right\}$
$+\delta(x^{\prime}_{3}-x_{3})\left[\frac{x_{2}}{x^{\prime}_{2}}\theta(x^{\prime}_{2}-x_{2})+\frac{x_{1}}{x^{\prime}_{1}}\theta(x^{\prime}_{1}-x_{1})\right]\frac{\Gamma_{220}}{1-x^{\prime}_{3}}+\delta(x^{\prime}_{2}-x_{2})\left[\frac{x_{3}}{x^{\prime}_{3}}\theta(x^{\prime}_{3}-x_{3})+\frac{x_{1}}{x^{\prime}_{1}}\theta(x^{\prime}_{1}-x_{1})\right]\frac{\Gamma_{202}}{1-x^{\prime}_{2}}$
$+\delta(x^{\prime}_{1}-x_{1})\left[\frac{x_{2}}{x^{\prime}_{2}}\theta(x^{\prime}_{2}-x_{2})+\frac{x_{3}}{x^{\prime}_{3}}\theta(x^{\prime}_{3}-x_{3})\right]\frac{\Gamma_{022}}{1-x^{\prime}_{1}}\Bigg{\}}+{\cal O}(\alpha_{s}^{2}).$ (29)
By projecting the renormalized three-quark operator onto $\Phi_{3}$ using (3), one readily reproduce the ERBL evolution kernel of $\Phi_{3}(x)$, whose explicit expression can be found in Eq. (4.18) of Passek-Kumericki:2008uqr .
Appendix B Analytical expressions for the convolution of the hard-scattering kernel
In this appendix, we enumerate the analytical expressions of various short-distance coefficients defined in (22), whose numerical values are tabulated in Table 1.
Through NLO accuracy, we have
${\cal H}^{ijkl}$ $=$ $\frac{1600\pi^{2}\alpha_{s}^{2}}{3}\Big{\{}e_{u}\,c^{ijkl}_{0}\left[1+{\alpha_{s}\over\pi}\left(c^{ijkl}_{1}L_{\mu}+c^{ijkl}_{2}\right)\right]$ (30)
$+e_{d}\,d^{ijkl}_{0}\left[1+{\alpha_{s}\over\pi}(d^{ijkl}_{1}L_{\mu}+d^{ijkl}_{2})\right]+{\cal O}(\alpha_{s}^{2})\Big{\}}.$
The various entries read
  • ${\cal H}^{0000}$
    $c^{0000}_{0}=1,\quad d^{0000}_{0}=2,\quad c^{0000}_{1}=d_{1}=\frac{29}{6}$
    $c_{2}^{0000}=-92\zeta_{5}+\frac{3719}{30}\zeta_{3}-\frac{1897}{60},$
    $d_{2}^{0000}=\frac{3}{2}\zeta_{5}+\frac{209}{120}\zeta_{3}+\frac{1741}{120}.$ (31)
  • ${\cal H}^{1000}$
    $c^{1000}_{0}=\frac{49}{6},d^{1000}_{0}=-\frac{49}{6},\quad c^{1000}_{1}=d_{1}=\frac{97}{18}$
    $c_{2}^{1000}=d_{2}^{1000}=-100\zeta_{5}+\frac{3203}{40}\zeta_{3}+\frac{981829}{30240}.$ (32)
  • ${\cal H}^{1010}$
    $c_{0}^{1010}=\frac{1519}{9},\quad d_{0}^{1010}=\frac{1568}{9},\quad c^{1010}_{1}=d^{1010}_{1}=\frac{107}{18}$
    $c_{2}^{1010}=-\frac{55926\zeta_{5}}{217}+\frac{151409\zeta_{3}}{868}+\frac{11652817}{133920},$
    $d_{2}^{1010}=-\frac{22419\zeta_{5}}{224}+\frac{160177\zeta_{3}}{2240}+\frac{5656271}{120960}.$ (33)
  • ${\cal H}^{1100}$
    $c_{0}^{1100}=\frac{7}{2},\quad d_{0}^{1100}=-\frac{7}{2},\quad c^{1100}_{1}=d^{1100}_{1}=\frac{11}{2}$
    $c_{2}^{1100}=-\frac{130}{3}\zeta_{5}+\frac{5161}{72}\zeta_{3}-\frac{369}{32},$
    $d_{2}^{1100}=\frac{92}{3}\zeta_{5}-\frac{11953}{360}\zeta_{3}+\frac{38417}{1440}.$ (34)
  • ${\cal H}^{1110}$
    $c_{0}^{1110}=\frac{343}{9},\quad d_{0}^{1110}=-\frac{343}{9},\quad c^{1110}_{1}=d^{1110}_{1}=\frac{109}{18}$
    $c_{2}^{1110}=d_{2}^{1110}=-\frac{2400\zeta_{5}}{7}+\frac{447161\zeta_{3}}{1960}+\frac{23648419}{211680}.$
  • ${\cal H}^{1111}$
    $c_{0}^{1111}=\frac{539}{9},\quad d_{0}^{1111}=\frac{196}{9},\quad c^{1111}_{1}=d^{1111}_{1}=\frac{37}{6}$
    $c_{2}^{1111}=\frac{762}{11}\zeta_{5}-\frac{19183}{770}\zeta_{3}-\frac{4745929}{332640},$
    $d_{2}^{1111}=-\frac{141}{4}\zeta_{5}+\frac{17807}{560}\zeta_{3}+\frac{750391}{30240}.$ (36)
  • ${\cal H}^{2000}$
    $c_{0}^{2000}=\frac{427}{20},\quad d_{0}^{2000}=\frac{161}{20},\quad c^{2000}_{1}=d^{2000}_{1}=\frac{103}{18}$
    $c_{2}^{2000}=-\frac{74566}{427}\zeta_{5}+\frac{18937349}{153720}\zeta_{3}+\frac{110572513}{1844640},$
    $d_{2}^{2000}=\frac{1780}{161}\zeta_{5}-\frac{244199}{57960}\zeta_{3}+\frac{12993881}{695520}.$ (37)
  • ${\cal H}^{2010}$
    $c_{0}^{2010}=\frac{539}{6},\quad d_{0}^{2010}=-\frac{539}{6},\quad c^{2010}_{1}=d^{2010}_{1}=\frac{113}{18}$
    $c_{2}^{2010}=d_{2}^{2010}=-\frac{119124}{385}\zeta_{5}+\frac{1590773}{7700}\zeta_{3}+\frac{4637387}{44352}.$
  • ${\cal H}^{2011}$
    $c_{0}^{2011}=\frac{931}{10},\quad d_{0}^{2011}=\frac{49}{5},\quad c^{2011}_{1}=d^{2011}_{1}=\frac{115}{18}$
    $c_{2}^{2011}=-\frac{11792}{19}\zeta_{5}+\frac{19751197}{47880}\zeta_{3}+\frac{622684033}{3447360},$
    $d_{2}^{2011}=-\frac{7471}{14}\zeta_{5}+\frac{367657}{1008}\zeta_{3}+\frac{55780799}{362880}.$ (39)
  • ${\cal H}^{2020}$
    $c_{0}^{2020}=\frac{23471}{100},\quad d_{0}^{2020}=\frac{1372}{25},\quad c^{2020}_{1}=d^{2020}_{1}=\frac{119}{18}$
    $c_{2}^{2020}=-\frac{3537990}{3353}\zeta_{5}+\frac{46755193}{67060}\zeta_{3}+\frac{702816409}{2414160},$
    $d_{2}^{2020}=\frac{430209}{784}\zeta_{5}-\frac{2824781}{7840}\zeta_{3}-\frac{16409957}{161280}.$ (40)
  • ${\cal H}^{2100}$
    $c_{0}^{2100}=\frac{35}{4},\quad d_{0}^{2100}=-\frac{35}{4},\quad c^{2100}_{1}=d^{2100}_{1}=\frac{107}{18}$
    $c_{2}^{2100}=d_{2}^{2100}=\frac{2732}{35}\zeta_{5}-\frac{179843}{4200}\zeta_{3}-\frac{7131}{5600}.$ (41)
  • ${\cal H}^{2110}$
    $c_{0}^{2110}=d_{0}^{2110}=\frac{637}{4},\quad c^{2110}_{1}=d^{2110}_{1}=\frac{13}{2}$
    $c_{2}^{2110}=\frac{4908}{13}\zeta_{5}-\frac{447763}{1820}\zeta_{3}-\frac{49062911}{786240},$
    $d_{2}^{2110}=-\frac{9546}{91}\zeta_{5}+\frac{129473}{1820}\zeta_{3}+\frac{6383213}{112320}.$ (42)
  • ${\cal H}^{2111}$
    $c_{0}^{2111}=\frac{539}{12},\quad d_{0}^{2111}=-\frac{539}{12},\quad c^{2111}_{1}=d^{2111}_{1}=\frac{119}{18}$
    $c_{2}^{2111}=d_{2}^{2111}=\frac{47940}{77}\zeta_{5}-\frac{307441}{770}\zeta_{3}-\frac{88027997}{665280}.$
  • ${\cal H}^{2120}$
    $c_{0}^{2120}=\frac{3381}{40},\quad d_{0}^{2120}=-\frac{3381}{40},\quad c^{2120}_{1}=d^{2120}_{1}=\frac{41}{6}$
    $c_{2}^{2120}=d_{2}^{2120}=\frac{36348}{161}\zeta_{5}-\frac{1422557}{9660}\zeta_{3}-\frac{85715459}{4173120}.$
  • ${\cal H}^{2121}$
    $c_{0}^{2121}=d_{0}^{2121}=\frac{637}{4},\quad c^{2121}_{1}=d^{2121}_{1}=\frac{127}{18}$
    $c_{2}^{2121}=-\frac{94014}{91}\zeta_{5}+\frac{124647}{182}\zeta_{3}+\frac{675228271}{2358720},$
    $d_{2}^{2121}=\frac{57285}{91}\zeta_{5}-\frac{106337}{260}\zeta_{3}-\frac{32207713}{262080}.$ (45)
  • ${\cal H}^{2200}$
    $c_{0}^{2200}=\frac{16}{5},\quad d_{0}^{2200}=\frac{1}{10},\quad c^{2200}_{1}=d^{2200}_{1}=6$
    $c_{2}^{2200}=-\frac{3543}{16}\zeta_{5}+\frac{1924849}{11520}\zeta_{3}+\frac{7842193}{138240},$
    $d_{2}^{2200}=840\zeta_{5}-\frac{144079}{360}\zeta_{3}-\frac{1479289}{4320}.$ (46)
  • ${\cal H}^{2210}$
    $c_{0}^{2210}=\frac{49}{4},\quad d_{0}^{2210}=-\frac{49}{4},\quad c^{2210}_{1}=d^{2210}_{1}=\frac{59}{9}$
    $c_{2}^{2210}=d_{2}^{2210}=\frac{34068}{35}\zeta_{5}-\frac{109994}{175}\zeta_{3}-\frac{4467161}{20160}.$
  • ${\cal H}^{2211}$
    $c_{0}^{2211}=\frac{427}{60},\quad d_{0}^{2211}=\frac{161}{60},\quad c^{2211}_{1}=d^{2211}_{1}=\frac{20}{3}$
    $c_{2}^{2211}=-\frac{180108}{61}\zeta_{5}+\frac{1752946}{915}\zeta_{3}+\frac{140407409}{175680},$
    $d_{2}^{2211}=-\frac{8352}{23}\zeta_{5}+\frac{15283}{69}\zeta_{3}+\frac{9688567}{66240}.$ (48)
  • ${\cal H}^{2220}$
    $c_{0}^{2220}=\frac{6069}{200},\quad d_{0}^{2220}=\frac{987}{200},\quad c^{2220}_{1}=d^{2220}_{1}=\frac{62}{9}$
    $c_{2}^{2220}=\frac{4267860}{2023}\zeta_{5}-\frac{28050761}{20230}\zeta_{3}-\frac{25391852771}{52436160},$
    $d_{2}^{2220}=\frac{1066224}{329}\zeta_{5}-\frac{7005001}{3290}\zeta_{3}-\frac{6511339837}{8527680}.$ (49)
  • ${\cal H}^{2221}$
    $c_{0}^{2221}=\frac{287}{40},\quad d_{0}^{2221}=-\frac{287}{40},\quad c^{2221}_{1}=d^{2221}_{1}=\frac{64}{9}$
    $c_{2}^{2221}=d_{2}^{2221}=-\frac{2540124}{287}\zeta_{5}+\frac{16718021}{2870}\zeta_{3}+\frac{5492284859}{2479680}.$
  • ${\cal H}^{2222}$
    $c_{0}^{2222}=\frac{779}{100},\quad d_{0}^{2222}=\frac{43}{25},\quad c^{2222}_{1}=d^{2222}_{1}=\frac{43}{6}$
    $c_{2}^{2222}=\frac{3727080}{779}\zeta_{5}-\frac{12158634}{3895}\zeta_{3}-\frac{11846765099}{10095840},$
    $d_{2}^{2222}=-\frac{20277}{43}\zeta_{5}+\frac{571137}{1720}\zeta_{3}+\frac{278914037}{2229120}.$ (51)